Using the financial statements for the Snider Corporation, calculate the 13 basic ratios found in the chapter. SNIDER CORPORATION Balance Sheet December 31, 20X1 Assets Current assets Cash Marketable securities Accounts receivable (net) Inventory $ 54,200 28,800 180,000 244,000 $507,000 60,400 Total current assetS Investments $658,000 255,000 Plant and equipment. Less: Accumulated depreciation Net plant and equipment 403,000 Total assets $ 970,400 Liabilities and Stockholders Equity Current liabilities Accounts payable Notes payable Accrued taxes $ 94,800 78,400 14,500 $187,700 Total current liabilities Long-term liabilities Bonds payable Total liabilities 159,800 $ 347,500 Stockholders equity Preferred stock, $50 par value Common stock, $1 par value Capital paid in excess of par Retained earnings $ 100,000 80,000 190,000 252,900 $ 622,900 $ 970,400 Total stockholders equity Total liabilities and stockholders equitySNIDER CORPORATION Income Statement For the Year Ending December 31, 20X1 Sales (on credit) Cost of goods sold Gross profit Selling and administrative expenses Operating profit (EBIT) Interest expense Earnings before taxes (EBT) Taxes $2,070,000 1,377,000 $693,000 505,000* $ 188,000 34,200 $153,800 85,800 $ 68,000 Earnings after taxes (EAT) Includes $35,800 in lease payments Using the above financial statements for the Snider Corporation, calculate the following ratios a. Profitability ratios. (Do not round intermediate calculations. Input your answers as a percent rounded to 2 decimal places.) Profitability Ratios Profit margin Return on assets (investment) Return on equityb. Assets utilization ratios. (Do not round intermediate calculations. Round your answers to 2 decimal places.) Assets Utilization Ratios Receivable turnover Average collection Inventory turnover Fixed asset turnover Total asset turnover times days imes imes times c.Liquidity ratios. (Do not round intermediate calculations. Round your answers to 2 decimal places.) Liquidity Ratios Current ratio times Quick ratio times d. Debt utilization ratios. (Do not round intermediate calculations. Input your debt to total assets answer as a percent rounded to 2 decimal places. Round your other answers to 2 decimal places.) Debt Utilization Ratios Debt to total assets Times interest earned Fixed charge times times
Answer #1
Profit margin Net income/Sales 3.29% =68000/2070000
Return on assets Net income/Total assets 7.01% =68000/970400
Return on equity Net income/Total equity 10.92% =68000/622900
Receivable turnover Sales/Accounts receivable    11.50 =2070000/180000
Average collection period Accounts receivable/Sales per day    31.30 =180000/(2070000/360)
Inventory turnover Cost of goods sold/Inventory      5.64 =1377000/244000
Fixed asset turnover Sales/Fixed assets      5.14 =2070000/403000
Total asset turnover Sales/Total assets      2.13 =2070000/970400
Current ratio Current assets/Current liabilities      2.70 =507000/187700
Quick ratio (Current assets-Inventory)/Current liabilities      1.40 =(507000-244000)/187700
Debt to total assets Total liabilities/Total assets      0.36 =347500/970400
Times interest earned EBIT/Interest expense      5.50 =188000/34200
Fixed charge coverage (EBIT+Fixed charges)/(Interest expense+Fixed charges)      3.20 =(188000+35800)/(34200+35800)